# Optimal Low-Depth Quantum Signal-Processing Phase Estimation

### v1

@@ D.4 Peak regression and peak fitting
When the number of additional samples $n$ is large enough, the estimator derived from the minimization problem is expected to be unbiased and the variance scales as $\mathcal{O}(1/(d^{2}nM))$ according to the M-estimation theory [13]. However, the practical implementation of these estimators is easily affected by the complex landscape of the minimization problem. Meanwhile, the sub-optimality and the run time of black-box optimization algorithms also limits the use of these estimator.
To overcome the difficulty due to the complex landscape of nonlinear regression, we propose another technique to improve the accuracy of the swap-angle estimator by fitting the peak of the amplitude function $\mathfrak{f}_{d}(\omega-\varphi,\theta)$. We observe that the amplitude function is well captured by a parabola on the interval $\mathcal{I}:=\left[\varphi-\frac{\pi}{2d},\varphi+\frac{\pi}{2d}\right]$. Consider $n_{\mathrm{pf}}$ equally spaced sample points on the interval $\mathcal{I}$: $\omega_{j}^{\mathrm{(pf)}}=\hat{\varphi}_{\mathrm{pri}}+\frac{\pi}{d}\left(\frac{j}{n_{\mathrm{pf}}-1}-\frac{1}{2}\right)$ where $j=0,1,\cdots,n_{\mathrm{pf}}-1$. We find the best parabola fitting the sampled data $\mathfrak{f}_{d}^{\mathrm{exp}}\left(\omega_{j}^{\mathrm{(pf)}}\right)$ whose maximum $\mathfrak{f}_{d}^{\mathrm{(pf\ max)}}$ attains at $\omega^{\mathrm{(pf\ max)}}$. Given that $\hat{\varphi}_{\mathrm{pri}}$ is an accurate estimator of the angle $\varphi$, we accept the parabolic fitting result if the peak location does not deviate $\hat{\varphi}_{\mathrm{pri}}$ beyond some threshold $\varepsilon^{\mathrm{thr}}$, namely, the fitting is accepted if |$\absolutevalue{\omega^{\mathrm{(pf\ max)}}-\hat{\varphi}_{\mathrm{pri}}}<\varepsilon^{\mathrm{thr}}$. Upon the acceptance, the estimator is $\hat{\theta}_{\mathrm{pf}}:=\mathfrak{f}_{d}^{\mathrm{(pf\ max)}}/d$. Ignoring the systematic bias caused by the overshooting of $\hat{\varphi}\neq\varphi$, the variance of the estimator is approximately $\mathcal{O}\left(\frac{1}{d^{2}n_{\mathrm{pf}}}\right)$. The detailed procedure is given in Section D.4.
@@
  Initiate real-valued data vectors $\vec{\mathfrak{p}}^{\mathrm{exp}},\vec{\mathfrak{w}}\in\mathbb{R}^{n}$.
@@
     Peform the quantum circuit in Figure 1 (or Figure 8) and measure the transition probabilities $p_{X}^{\mathrm{exp}}(\omega_{j})$ and $p_{Y}^{\mathrm{exp}}(\omega_{j})$.
     Set $\vec{\mathfrak{p}}^{\mathrm{exp}}_{j}\leftarrow\sqrt{\left(p_{X}^{\mathrm{exp}}(\omega_{j})-\frac{1}{2}\right)^{2}+\left(p_{Y}^{\mathrm{exp}}(\omega_{j})-\frac{1}{2}\right)^{2}}$ and $\vec{\mathfrak{w}}_{j}\leftarrow\omega_{j}$.
@@
  Fit $\left(\vec{\mathfrak{w}},\vec{\mathfrak{p}}^{\mathrm{exp}}\right)$ with respect to to parabolic model $\mathfrak{p}=\beta_{0}\left(\mathfrak{w}-\beta_{1}\right)^{2}+\beta_{2}$.
@@
D.5 Numerical performance of QSPE on FsimGate against Monte Carlo sampling error
@@
To numerically test the performance of QSPE and validate the analysis in the presence of Monte Carlo sampling error, we simulate the quantum circuit and perform the inference. In Figure 10, we plot the squared error of each estimator as a function of the number of FsimGates $d$ in each quantum circuit. Consequentially, each data point is the mean squared error (MSE), which is a metric of the performance according to the bias-variance decomposition $\mathrm{MSE}=\mathrm{Var}+\mathrm{bias}^{2}$. The numerical results in Figure 10 indicates that although $\theta=1\times 10^{-3}$ is small, QSPE estimators achieve an accurate estimation with a very small $d$. The numerical results also show that the performance of the estimator does not significantly depend on the value of the single-qubit phase $\varphi$. Meanwhile, using the peak fitting in Section D.4, the variance in $\theta$-estimation is improved so that the MSE curve is lowered. Zooming the MSE curve in log-log scale, the curve scales as a function of $d$ as the theoretically derived variance scaling in Theorem 2. We will discuss the scaling of the variance in Appendix F in more details.
@@
In Figure 11, we perform the numerical simulation with variable swap angle $\theta$ and number of measurement samples $M$. The numerical results show that the accuracy of $\varphi$-estimation is more vulnerable to decreasing $\theta$. This is explainable from the theoretically derived variance in Theorem 2 which depends on the swap angle as $1/\theta^{2}$. Although the theoretical variance of $\theta$ is expected to be invariant for different $\theta$ values, the numerical results show that the MSE of $\theta$-estimation gets larger when smaller $\theta$ is used, and the scaling of the curve differs from the classical scaling $1/M$. The reason is that when $\theta\leq 5\times 10^{-4}$, the SNR is not large enough so that the theoretical derivation can be justified. When using a bigger $d$ or $M$, the curve will converge to the theoretical derivation. When $\theta=1\times 10^{-3}$, the setting of the experiments is enough to get a large enough SNR. Hence, the scaling of the MSE curves in the bottom panels in Figure 11 agrees with the classical scaling $1/M$ of Monte Carlo sampling error.
@@
Figure 10: Squared error of estimators as a function of the number of FsimGate’s. The only source of noise in the numerical experiments is Monte Carlo sampling error. The number of measurement samples is $M=1\times 10^{5}$, and $n_{\mathrm{pf}}=15$ is used in the peak fitting. The swap angle is set to $\theta=1\times 10^{-3}$ and the phase parameter is set to $\chi=5\pi/32$. The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
@@
Figure 11: Squared error of estimators as a function of the number of measurement samples. The only source of noise in the numerical experiments is Monte Carlo sampling error. The circuit degree is set to $d=50$ and the FsimGate phase parameter is set to $\chi=5\pi/32$. $n_{\mathrm{pf}}=15$ is used in the peak fitting. The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
@@
Appendix E Solving QSPE with arbitrary swap-angle value
@@
In the early section, we derive robust statistical estimators when $d\theta\leq 1$. These estimators are based on linear statistical models which relies on the approximation of Fourier coefficients in the desired regime. In this section, we aim to generalize the solution to arbitrary swap-angle value which expands the use of QSPE to generic $U$-gates.
@@
Recall that in Theorem 9, we show that by performing Fourier transformation, the swap angle and phase angles are fully decoupled in terms of the dependencies in amplitude and phase. Furthermore, the analysis in Appendix D reveals that the noise magnitude, namely, the variance of the noise, in the Fourier space is reduced by a factor of $d$. Consequently, these results suggest the design of algorithms in Fourier space.
@@
According to Theorems 9 and 39, the amplitudes of Fourier coefficients are exactly computable by solving the integral:
@@
    Nonconvexity. Each function $A_{k}$ is nonconvex which contributes to the difficulty in solving nonlinear equation. Furthermore, the function is not injective which makes $A_{k}^{-1}$ multi-valued. These render the numerical solution to the system of nonlinear equations challening.
@@
This procedure is depicted in Figure 12 where each panel stands for an individual nonlinear equation with two distinct choice of $\alpha_{k}=\pm 1$. In each panel, all intersection points form the potential solution set $A_{k}^{-1}(\alpha_{k}\cdot\absolutevalue{c_{k}^{\mathrm{exp}}})$. As the degree parameter is set to $d=5$, the system has nine nonlinear equations. By solving them using the previously discussed interval-based method, we obtain the counter which is visualized as a histogram in Figure 13. We see the full satisfaction is attained at $\theta=1$ and $\theta=\pi-1$. They form the final solution output of the algorithm. It is worth noting that these two $\theta$-values are equivalent as the distinction in negating cosine can be accounted by redefining other phase angles.
To estimate phase angle $\varphi$, we note that Theorem 9 indicates that the $\varphi$-estimation procedure in general case is identical to that when $\theta$ is small. Hence, we may still use the derived estimator to estimate $\varphi$ by taking the sequential phase difference in the phases of Fourier coefficients. However, we note that the $\theta$-value may affect the estimation accuracy of $\varphi$ as the value of $A_{k}(\theta)$ modulate the magnitude of the Fourier mode and limit the signal-to-noise ratio as revealed in the analysis in Appendix D.
In Figure 14, we test the algorithm performance with variable $\theta$-values. We see that the absolute error in $\theta$-estimation remains well bounded below $5\times 10^{-4}$ except for the singularity near $\pi/2$. At this value, the variable part of the transition probabilities vanishes and $\mathfrak{h}=0$. Hence, angle inference becomes increasingly challenging due to the absence of information. It is also worth noting that as $\theta$ gets close to $\pi/2$, the magnitude of $\cos(\theta)$ becomes more vanishing. Hence, the signal strength of the phase angle dependent part is increasingly weakened comparing to noise. This low signal-to-noise ratio leads to the increase in the $\varphi$-estimation error in Figure 14.
@@
  Initiate a complex-valued data vector $\vec{\mathfrak{h}}^{\mathrm{exp}}\in\mathbb{C}^{2d-1}$.
@@
     Perform the quantum circuit in Figure 1 (or Figure 8) and measure the transition probabilities $p_{X}^{\mathrm{exp}}(\omega_{j})$ and $p_{Y}^{\mathrm{exp}}(\omega_{j})$.
     Set $\vec{\mathfrak{h}}^{\mathrm{exp}}_{j}\leftarrow p_{X}^{\mathrm{exp}}(\omega_{j})-\frac{1}{2}+\mathrm{i}\left(p_{Y}^{\mathrm{exp}}(\omega_{j})-\frac{1}{2}\right)$.
@@
  Compute the Fourier coefficients $\vec{c}^{\mathrm{exp}}=\mathsf{FFT}\left(\vec{\mathfrak{h}}^{\mathrm{exp}}\right)$.
  Compute estimates $\hat{\varphi}$ according to Definition 17 using $\mathsf{phase}(\vec{c}^{\mathrm{exp}})$.
@@
Figure 12: An example of formulating QSPE as solving systems of nonlinear equations. We set $d=5$ and $M=1\times 10^{5}$. The relevant angles are set to $\theta=1,\varphi=\pi/16$. In each panel, the solid curve is the amplitude of the Fourier coefficients derived by solving the defining integral. The two horizontal dashed lines stand for the experimental amplitude with positive or negative signs ($\alpha_{k}=\pm 1$). The intersecting points are candidate solutions $A_{k}^{-1}(\alpha_{k}\cdot\absolutevalue{c_{k}^{\mathrm{exp}}})$.
Figure 13: An example of the counter of interval-based algorithm for solving QSPE. The setup is identical to that in Figure 12. The $\theta$-values satisfy all nine nonlinear equations are the output of the algorithm, which are estimators of the swap angle.
Figure 14: Estimation errors of $\theta$ and $\varphi$ using the generalized QSPE estimation algorithm for general swap angles. We set $d=5$ and $M=1\times 10^{5}$. We fix $\varphi=\pi/16$ and make $\theta$ variable. The shaded area stands for the confidence interval (error bar) determined from ten independent repetitions.
Appendix F Lower bounding the performance of QSPE
@@
when $d$ is large enough. However, theoretical analysis in Theorem 2 and numerical simulation in Figure 10 show that the variance of the $\varphi$-estimator in QSPE depends on the parameter $d$ as $\mathrm{Var}\left(\hat{\varphi}\right)\sim 1/d^{4}$. In this section, we will analyze this nontrivial counterintuitive result. In the end, we prove that for a fixed unknown FsimGate, the $1/d^{4}$-dependency only appears in the pre-asymptotic regime where the condition of the theorems holds, i.e., $d\theta\ll 1$. When passing to the limit of large enough $d$, the variances of statistical estimators agree with that suggested by the Heisenberg limit. Although such faster than Heisenberg limit scaling only applies in a finite range of circuit depth ($d\theta\ll 1$), it has drastically increased our metrology performance in practice against time-dependent errors, and thus deserves further investigation in its generalization to other domains of noise learning.
F.1 Pre-asymptotic regime $d\ll 1/\theta$
@@
Remarkably, the variances of QSPE estimators in Theorem 2 exactly match the optimality given in Equations 143 and 144. We thus proves the optimality of our QSPE estimator for inferring gate parameter $\theta$ and $\varphi$. Moreover, we like to point out that the faster than Heisenberg-limit scaling of parameter $\varphi$ in this asymptotic regime is critical to the successful experimental deployment of our methods. This is because the dominant time-dependent error results in a time-dependent drift error in $\varphi$, and a faster convergence in circuit depth provides faster metrology runtime to minimize such drift error during the measurements.
F.2 Asymptotic regime $d\to\infty$
@@
F.3 Numerical results
@@
where the Fisher information matrix is element-wisely defined in Equation 136. At the same time, we also compute the approximation to the optimal variance in the pre-asymptotic regime $d\ll 1/\theta$ derived in Equations 143, 144 and 145. The numerical results are given in Figure 3(b). It can be seen that the approximated optimal variance agrees very well with the exact CRLB. In the asymptotic regime with large enough $d$, the optimal variance scaling given by the CRLB is as predicted in Equation 147. Furthermore, the numerical results validate that there exists a nontrivial transition around $d\approx 1/\theta$ making the optimal variance scalings completely different in the pre-asymptotic and asymptotic regime.
@@
Figure 15: Variance and mean-square error (MSE) of QSPE estimators. The left panel corresponds to the case where $\theta$ is relatively large and $d\theta\ll 1$ condition fails quickly at around $d=10$, beyond which bias dominates the estimator’s MSE since our inference model assumption ($d\theta\ll 1$) fails. The right panel corresponds to the case where $d\theta\ll 1$ condition holds all the way to around $d=100$. The single-qubit phases are set to $\varphi=\pi/16$ and $\chi=5\pi/32$. The number of measurement samples is set to $M=1\times 10^{5}$. Each data point is derived from $100$ independent repetitions.
@@
To prove the optimality of QSPE and investigate the situation where the conditions for deriving QSPE hold, we numerically estimate the variances of QSPE estimators and compare them with the derived optimal variances in the pre-asymptotic regime in Equations 143 and 144. The QSPE estimators are derived by approximating the original statistical inference problem by a linear model. When $d$ gets large, the model violation due to the approximation contributes to the bias of QSPE estimators. We compute the mean-square error (MSE) and using the bias-variance decomposition $\mathrm{MSE}=\mathrm{Var}+\mathrm{bias}^{2}$ to quantify the bias. The numerical results are displayed in Figure 15. Our simulation shows that the bias of $\theta$-estimator dominates the MSE and contaminates the inference accuracy after $d$ becomes larger than a threshold determined by the pre-asymptotic regime $d\theta\ll 1$. Despite the bias due to the model violation, the MSE of the $\theta$-estimator still achieves some accuracy of order $\theta^{2}$ which suggests that the $\theta$-estimator might give a reasonable estimation of a similar order with model violation in larger $d$. The numerical results show that the $\varphi$-estimator is more robust where the MSE deviates significantly from the theoretical scaling in the pre-asymptotic regime after $d\geq 1/\theta$ is large enough to pass to the asymptotic regime. Furthermore, the MSE well matches the variance which implies that the bias in $\varphi$-estimator is always small. The difference in the robustness of the $\theta$- and $\varphi$-estimators is credited to the construction of QSPE in which the inferences of $\theta$ and $\varphi$ are completely decoupled due to the data post-processing using FFT.
@@
Figures 3(b) and 15 suggest the following. (1) In the pre-asymptotic regime, QSPE estimators achieve the optimality in the sense of saturating the Cramér-Rao lower bound and exhibit robustness against time-dependent errors in $\varphi$ in both simulation and experimental deployments. Furthermore, the construction of QSPE estimators only involves direct algebraic operations rather than iterative optimization, and the reduced inference problems in Fourier space are linear statistical models whose global optimum is unique for each realization. This not only enables the fast and reliable data post-processing but also allows us to analyze its performance analytically. (2) Passing to the asymptotic regime, given the significant bias of $\theta$-estimator and the sharp transition of the variance of $\varphi$-estimator, one has to use other estimators to saturate the optimal variance scaling and unbiasedness, for example, maximum-likelihood estimators (MLE). Furthermore, the analysis based on the Cramér-Rao lower bound is made by fixing the data generation (measuring quantum circuits) but varying data post-processing.
@@
F.4 Comparing with quantum Cramer-Rao lower bound

### v2

@@ Acknowledgments
When the number of additional samples $n$ is large enough, the estimator derived from the minimization problem is expected to be unbiased and the variance scales as $\mathcal{O}(1/(d^{2}nM))$ according to the M-estimation theory [15]. However, the practical implementation of these estimators is easily affected by the complex landscape of the minimization problem. Meanwhile, the sub-optimality and the run time of black-box optimization algorithms also limit the use of these estimators.
To overcome the difficulty due to the complex landscape of nonlinear regression, we propose another technique to improve the accuracy of the swap-angle estimator by fitting the peak of the amplitude function $\mathfrak{f}_{d}(\omega-\varphi,\theta)$. We observe that the amplitude function is well captured by a parabola on the interval $\mathcal{I}:=\left[\varphi-\frac{\pi}{2d},\varphi+\frac{\pi}{2d}\right]$. Consider $n_{\mathrm{pf}}$ equally spaced sample points on the interval $\mathcal{I}$: $\omega_{j}^{\mathrm{(pf)}}=\hat{\varphi}_{\mathrm{pri}}+\frac{\pi}{d}\left(\frac{j}{n_{\mathrm{pf}}-1}-\frac{1}{2}\right)$ where $j=0,1,\cdots,n_{\mathrm{pf}}-1$. We find the best parabola fitting the sampled data $\mathfrak{f}_{d}^{\mathrm{exp}}\left(\omega_{j}^{\mathrm{(pf)}}\right)$ whose maximum $\mathfrak{f}_{d}^{\mathrm{(pf\ max)}}$ attains at $\omega^{\mathrm{(pf\ max)}}$. Given that $\hat{\varphi}_{\mathrm{pri}}$ is an accurate estimator of the angle $\varphi$, we accept the parabolic fitting result if the peak location does not deviate $\hat{\varphi}_{\mathrm{pri}}$ beyond some threshold $\varepsilon^{\mathrm{thr}}$, namely, the fitting is accepted if |$\absolutevalue{\omega^{\mathrm{(pf\ max)}}-\hat{\varphi}_{\mathrm{pri}}}<\varepsilon^{\mathrm{thr}}$. Upon the acceptance, the estimator is $\hat{\theta}_{\mathrm{pf}}:=\mathfrak{f}_{d}^{\mathrm{(pf\ max)}}/d$. Ignoring the systematic bias caused by the overshooting of $\hat{\varphi}\neq\varphi$, the variance of the estimator is approximately $\mathcal{O}\left(\frac{1}{d^{2}n_{\mathrm{pf}}}\right)$. The detailed procedure is given in Section Supplementary Information 4.4.
@@
  Initiate real-valued data vectors $\bm{\mathfrak{p}}^{\mathrm{exp}},\bm{\mathfrak{w}}\in\mathbb{R}^{n}$.
@@
     Peform the quantum circuit in Figure 1 in the main text (or Figure 3) and measure the transition probabilities $p_{X}^{\mathrm{exp}}(\omega_{j})$ and $p_{Y}^{\mathrm{exp}}(\omega_{j})$.
     Set $\bm{\mathfrak{p}}^{\mathrm{exp}}_{j}\leftarrow\sqrt{\left(p_{X}^{\mathrm{exp}}(\omega_{j})-\frac{1}{2}\right)^{2}+\left(p_{Y}^{\mathrm{exp}}(\omega_{j})-\frac{1}{2}\right)^{2}}$ and $\bm{\mathfrak{w}}_{j}\leftarrow\omega_{j}$.
@@
  Fit $\left(\bm{\mathfrak{w}},\bm{\mathfrak{p}}^{\mathrm{exp}}\right)$ with respect to to parabolic model $\mathfrak{p}=\beta_{0}\left(\mathfrak{w}-\beta_{1}\right)^{2}+\beta_{2}$.
@@
Supplementary Information 4.5 Numerical performance of QSPE on FsimGate against Monte Carlo sampling error
@@
To numerically test the performance of QSPE and validate the analysis in the presence of Monte Carlo sampling error, we simulate the quantum circuit and perform the inference. In Figure 5, we plot the squared error of each estimator as a function of the number of FsimGates $d$ in each quantum circuit. Consequentially, each data point is the mean squared error (MSE), which is a metric of the performance according to the bias-variance decomposition $\mathrm{MSE}=\mathrm{Var}+\mathrm{bias}^{2}$. The numerical results in Figure 5 indicate that although $\theta=1\times 10^{-3}$ is small, QSPE estimators achieve an accurate estimation with a very small $d$. The numerical results also show that the performance of the estimator does not significantly depend on the value of the single-qubit phase $\varphi$. Meanwhile, using the peak fitting in Section Supplementary Information 4.4, the variance in $\theta$-estimation is improved so that the MSE curve is lowered. Zooming the MSE curve in log-log scale, the curve scales as a function of $d$ as the theoretically derived variance scaling in Theorem 2 in the main text. We will discuss the scaling of the variance in Section Supplementary Information 6 in more detail.
@@
In Figure 6, we perform the numerical simulation with variable swap angle $\theta$ and the number of measurement samples $M$. The numerical results show that the accuracy of $\varphi$-estimation is more vulnerable to decreasing $\theta$. This is explainable from the theoretically derived variance in Theorem 2 in the main text which depends on the swap angle as $1/\theta^{2}$. Although the theoretical variance of $\theta$ is expected to be invariant for different $\theta$ values, the numerical results show that the MSE of $\theta$-estimation gets larger when smaller $\theta$ is used, and the scaling of the curve differs from the classical scaling $1/M$. The reason is that when $\theta\leq 5\times 10^{-4}$, the SNR is not large enough so that the theoretical derivation can be justified. When using a bigger $d$ or $M$, the curve will converge to the theoretical derivation. When $\theta=1\times 10^{-3}$, the setting of the experiments is enough to get a large enough SNR. Hence, the scaling of the MSE curves in the bottom panels in Figure 6 agrees with the classical scaling $1/M$ of Monte Carlo sampling error.
@@
Supplementary Figure 5: Squared error of estimators as a function of the number of FsimGate’s. The only source of noise in the numerical experiments is the Monte Carlo sampling error. The number of measurement samples is $M=1\times 10^{5}$, and $n_{\mathrm{pf}}=15$ is used in the peak fitting. The swap angle is set to $\theta=1\times 10^{-3}$ and the phase parameter is set to $\chi=5\pi/32$. The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
@@
Supplementary Figure 6: Squared error of estimators as a function of the number of measurement samples. The only source of noise in the numerical experiments is the Monte Carlo sampling error. The circuit degree is set to $d=50$ and the FsimGate phase parameter is set to $\chi=5\pi/32$. $n_{\mathrm{pf}}=15$ is used in the peak fitting. The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
@@
Supplementary Information 5 Solving QSPE with arbitrary swap-angle value
@@
In the early section, we derive robust statistical estimators when $d\theta\leq 1$. These estimators are based on linear statistical models which rely on the approximation of Fourier coefficients in the desired regime. In this section, we aim to generalize the solution to arbitrary swap-angle value which expands the use of QSPE to generic $U$-gates.
@@
Recall that in Theorem 7, we show that by performing Fourier transformation, the swap angle and phase angles are fully decoupled in terms of the dependencies in amplitude and phase. Furthermore, the analysis in Section Supplementary Information 4 reveals that the noise magnitude, namely, the variance of the noise, in the Fourier space is reduced by a factor of $d$. Consequently, these results suggest the design of algorithms in Fourier space.
@@
According to Theorems 7 and 27, the amplitudes of Fourier coefficients are exactly computable by solving the integral:
@@
    Nonconvexity. Each function $A_{k}$ is nonconvex which contributes to the difficulty in solving nonlinear equation. Furthermore, the function is not injective which makes $A_{k}^{-1}$ multi-valued. These render the numerical solution to the system of nonlinear equations challenging.
@@
This procedure is depicted in Figure 7 where each panel stands for an individual nonlinear equation with two distinct choices of $\alpha_{k}=\pm 1$. In each panel, all intersection points form the potential solution set $A_{k}^{-1}(\alpha_{k}\cdot\absolutevalue{c_{k}^{\mathrm{exp}}})$. As the degree parameter is set to $d=5$, the system has nine nonlinear equations. By solving them using the previously discussed interval-based method, we obtain the counter which is visualized as a histogram in Figure 8. We see the full satisfaction is attained at $\theta=1$ and $\theta=\pi-1$. They form the final solution output of the algorithm. It is worth noting that these two $\theta$-values are equivalent as the distinction in negating cosine can be accounted by redefining other phase angles.
To estimate phase angle $\varphi$, we note that Theorem 7 indicates that the $\varphi$-estimation procedure in the general case is identical to that when $\theta$ is small. Hence, we may still use the derived estimator to estimate $\varphi$ by taking the sequential phase difference in the phases of Fourier coefficients. However, we note that the $\theta$-value may affect the estimation accuracy of $\varphi$ as the value of $A_{k}(\theta)$ modulates the magnitude of the Fourier mode and limits the signal-to-noise ratio as revealed in the analysis in Section Supplementary Information 4.
In Figure 9, we test the algorithm performance with variable $\theta$-values. We see that the absolute error in $\theta$-estimation remains well bounded below $5\times 10^{-4}$ except for the singularity near $\pi/2$. At this value, the variable part of the transition probabilities vanishes and $\mathfrak{h}=0$. Hence, angle inference becomes increasingly challenging due to the absence of information. It is also worth noting that as $\theta$ gets close to $\pi/2$, the magnitude of $\cos(\theta)$ becomes more vanishing. Hence, the signal strength of the phase angle dependent part is increasingly weakened compared to noise. This low signal-to-noise ratio leads to the increase in the $\varphi$-estimation error in Figure 9.
@@
  Initiate a complex-valued data vector $\bm{\mathfrak{h}}^{\mathrm{exp}}\in\mathbb{C}^{2d-1}$.
@@
     Perform the quantum circuit in Figure 1 in the main text (or Figure 3) and measure the transition probabilities $p_{X}^{\mathrm{exp}}(\omega_{j})$ and $p_{Y}^{\mathrm{exp}}(\omega_{j})$.
     Set $\bm{\mathfrak{h}}^{\mathrm{exp}}_{j}\leftarrow p_{X}^{\mathrm{exp}}(\omega_{j})-\frac{1}{2}+\mathrm{i}\left(p_{Y}^{\mathrm{exp}}(\omega_{j})-\frac{1}{2}\right)$.
@@
  Compute the Fourier coefficients $\mathbf{c}^{\mathrm{exp}}=\mathsf{FFT}\left(\bm{\mathfrak{h}}^{\mathrm{exp}}\right)$.
  Compute estimates $\hat{\varphi}$ according to Definition 15 using $\mathsf{phase}(\mathbf{c}^{\mathrm{exp}})$.
@@
Supplementary Figure 7: An example of formulating QSPE as solving systems of nonlinear equations. We set $d=5$ and $M=1\times 10^{5}$. The relevant angles are set to $\theta=1,\varphi=\pi/16$. In each panel, the solid curve is the amplitude of the Fourier coefficients derived by solving the defining integral. The two horizontal dashed lines stand for the experimental amplitude with positive or negative signs ($\alpha_{k}=\pm 1$). The intersecting points are candidate solutions $A_{k}^{-1}(\alpha_{k}\cdot\absolutevalue{c_{k}^{\mathrm{exp}}})$.
Supplementary Figure 8: An example of the counter of interval-based algorithm for solving QSPE. The setup is identical to that in Figure 7. The $\theta$-values satisfy all nine nonlinear equations are the output of the algorithm, which are estimators of the swap angle.
Supplementary Figure 9: Estimation errors of $\theta$ and $\varphi$ using the generalized QSPE estimation algorithm for general swap angles. We set $d=5$ and $M=1\times 10^{5}$. We fix $\varphi=\pi/16$ and make $\theta$ variable. The shaded area stands for the confidence interval (error bar) determined from ten independent repetitions.
Supplementary Information 6 Lower bounding the performance of QSPE
@@
when $d$ is large enough. However, theoretical analysis in Theorem 2 in the main text and numerical simulation in Figure 5 show that the variance of the $\varphi$-estimator in QSPE depends on the parameter $d$ as $\mathrm{Var}\left(\hat{\varphi}\right)\sim 1/d^{4}$. In this section, we will analyze this nontrivial counterintuitive result. In the end, we prove that for a fixed unknown FsimGate, the $1/d^{4}$-dependency only appears in the pre-asymptotic regime where the condition of the theorems holds, i.e., $d\theta\ll 1$. When passing to the limit of large enough $d$, the variances of statistical estimators agree with that suggested by the Heisenberg limit. Although such faster than Heisenberg limit scaling only applies in a finite range of circuit depth ($d\theta\ll 1$), it has drastically increased our metrology performance in practice against time-dependent errors, and thus deserves further investigation in its generalization to other domains of noise learning.
Supplementary Information 6.1 Pre-asymptotic regime $d\ll 1/\theta$
@@
Remarkably, the variances of QSPE estimators in Theorem 2 in the main text exactly match the optimality given in Equations 131 and 132. We thus proves the optimality of our QSPE estimator for inferring gate parameter $\theta$ and $\varphi$. Moreover, we like to point out that the faster than Heisenberg-limit scaling of parameter $\varphi$ in this asymptotic regime is critical to the successful experimental deployment of our methods. This is because the dominant time-dependent error results in a time-dependent drift error in $\varphi$, and a faster convergence in circuit depth provides faster metrology runtime to minimize such drift error during the measurements.
Supplementary Information 6.2 Asymptotic regime $d\to\infty$
@@
Supplementary Information 6.3 Numerical results
@@
where the Fisher information matrix is element-wisely defined in Equation 124. At the same time, we also compute the approximation to the optimal variance in the pre-asymptotic regime $d\ll 1/\theta$ derived in Equations 131, 132 and 133. The numerical results are given in Figure 3 (b) in the main text. It can be seen that the approximated optimal variance agrees very well with the exact CRLB. In the asymptotic regime with large enough $d$, the optimal variance scaling given by the CRLB is as predicted in Equation 135. Furthermore, the numerical results validate that there exists a nontrivial transition around $d\approx 1/\theta$ making the optimal variance scalings completely different in the pre-asymptotic and asymptotic regime.
@@
Supplementary Figure 10: Variance and mean-square error (MSE) of QSPE estimators. The left panel corresponds to the case where $\theta$ is relatively large and $d\theta\ll 1$ condition fails quickly at around $d=10$, beyond which bias dominates the estimator’s MSE since our inference model assumption ($d\theta\ll 1$) fails. The right panel corresponds to the case where $d\theta\ll 1$ condition holds all the way to around $d=100$. The single-qubit phases are set to $\varphi=\pi/16$ and $\chi=5\pi/32$. The number of measurement samples is set to $M=1\times 10^{5}$. Each data point is derived from $100$ independent repetitions.
@@
To prove the optimality of QSPE and investigate the situation where the conditions for deriving QSPE hold, we numerically estimate the variances of QSPE estimators and compare them with the derived optimal variances in the pre-asymptotic regime in Equations 131 and 132. The QSPE estimators are derived by approximating the original statistical inference problem by a linear model. When $d$ gets large, the model violation due to the approximation contributes to the bias of QSPE estimators. We compute the mean-square error (MSE) and using the bias-variance decomposition $\mathrm{MSE}=\mathrm{Var}+\mathrm{bias}^{2}$ to quantify the bias. The numerical results are displayed in Figure 10. Our simulation shows that the bias of $\theta$-estimator dominates the MSE and contaminates the inference accuracy after $d$ becomes larger than a threshold determined by the pre-asymptotic regime $d\theta\ll 1$. Despite the bias due to the model violation, the MSE of the $\theta$-estimator still achieves some accuracy of order $\theta^{2}$ which suggests that the $\theta$-estimator might give a reasonable estimation of a similar order with model violation in larger $d$. The numerical results show that the $\varphi$-estimator is more robust where the MSE deviates significantly from the theoretical scaling in the pre-asymptotic regime after $d\geq 1/\theta$ is large enough to pass to the asymptotic regime. Furthermore, the MSE well matches the variance which implies that the bias in $\varphi$-estimator is always small. The difference in the robustness of the $\theta$- and $\varphi$-estimators is credited to the construction of QSPE in which the inferences of $\theta$ and $\varphi$ are completely decoupled due to the data post-processing using FFT.
@@
Figure 10 and Figure 3 (b) in the main text suggest the following. (1) In the pre-asymptotic regime, QSPE estimators achieve the optimality in the sense of saturating the Cramér-Rao lower bound and exhibit robustness against time-dependent errors in $\varphi$ in both simulation and experimental deployments. Furthermore, the construction of QSPE estimators only involves direct algebraic operations rather than iterative optimization, and the reduced inference problems in Fourier space are linear statistical models whose global optimum is unique for each realization. This not only enables the fast and reliable data post-processing but also allows us to analyze its performance analytically. (2) Passing to the asymptotic regime, given the significant bias of $\theta$-estimator and the sharp transition of the variance of $\varphi$-estimator, one has to use other estimators to saturate the optimal variance scaling and unbiasedness, for example, maximum-likelihood estimators (MLE). Furthermore, the analysis based on the Cramér-Rao lower bound is made by fixing the data generation (measuring quantum circuits) but varying data post-processing.
@@
Supplementary Information 6.4 Comparing with quantum Cramér-Rao lower bound